OpenStudy (anonymous):

4x – 4 + 3x = 3x + 5 – 5x

OpenStudy (anonymous):

4x + 3x - 3x + 5x = 4 + 5 9x = 9 x = 1

OpenStudy (anonymous):

How did you get 9x?

OpenStudy (anonymous):

4x+3x=7x 7x-3x=4x 4x+5x=9x

OpenStudy (anonymous):

Add all the liked terms (ones with x's next to them, together) and then add the ones with no x's together.